OpenStudy (anonymous):

If R1 and R2 are increasing at rates of 0.3 Ω/s and 0.2 Ω/s, respectively, how fast is R changing when R1 = 70 Ω and R2 = 110 Ω? (Round your answer to three decimal places.) 1/R=1/R1+1/R2

OpenStudy (anonymous):

They want you to use implicit differentiation.

OpenStudy (anonymous):

\[ \frac{d}{dt}\frac1R=-\frac{dR/dt}{R^2} \]

OpenStudy (anonymous):

So \[ -\frac{R'}{R^2}=-\left(\frac{R_1'}{R_1^2}+\frac{R_2'}{R_2^2}\right) \]

OpenStudy (anonymous):

R1 = 70 Ω and R2 = 110 Ω? R1 and R2 are increasing at rates of 0.3 Ω/s and 0.2 Ω/s,

OpenStudy (anonymous):

Plug them in.

OpenStudy (anonymous):

i got .000000042

OpenStudy (anonymous):

Okay, check in the back of the book?

OpenStudy (anonymous):

its online. nothing in the back of the book

OpenStudy (anonymous):

I get \(R=42.7\)

OpenStudy (anonymous):

\(R' \approx 0.142\)